OpenStudy (anonymous):

for a medal: how do i find the domain of cosx / (2+sinx)? i set the denominator equal to 0 and got sinx= -2. then i did sin inverse and got an error. what is the correct way to solve this?

OpenStudy (campbell_st):

well thats he correct method, the problem is value of sin(x) is always between -1 and 1 that is the range of the function is \[-1\le \sin(x) \le 1\] so you can't find an angle that will result in sin(x) = -2 because the angle doesn't exist so the domain is all real x. in other words, you can put any value of x in the denominator and you'll get a positive value.. I've attached a graph of the curve... the horizontal scale is in radians hope it helps